(1) calculate the equilibrium price of the following demand and supply equation

Qd=150-15/2p

Qs=10+5p

(2)The demand function for monopolist is given by p=85+11q-3q2.its cost function

C=30q-9q2+3q2+11 where q>0.find the price output and the profit of the monopolist


1
Expert's answer
2021-03-19T12:44:19-0400

Equilibrium price

(Qd = 120-5P) = (Qs = -30 + 10P)

1OP + 5P = 120 + 30

15P = 150

P = 150/15 =$10


Equilibrium Quantity (Qd)

Qd = 120-5P

Qd = 120-5(10)

Qd = 120-50

Qd = 70